HELP!!! i dont get this problem Please show your work!!!
t-2/t = 2/t+3

Found 2 solutions by solver91311, unreal4evr:
Answer by solver91311(24713)



Presuming your ambiguously rendered equation is actually:



Cross multiply:



Now all you have to do is FOIL the two binomials in the LHS, collect terms to put the equation into standard quadratic form, and solve the resulting quadratic by your choice of method.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by unreal4evr(11)

t - 2/t = 2/t + 3
t = 4/t + 3
t - 3 = 4/t
t2 - 3t = 4
t2 - 3t - 4 = 0


//You can solve this with the quadratic formula (a=1, b=-3, c=-4)
//Or you can rewrite the equation like so:


(t+1)(t-4) = 0


//Set each equal to 0


t+1 = 0
t = -1
t-4 = 0
t = 4


//Check
t = -1
(-1)-2/(-1) = 2/(-1) + 3
-1+2 = -2 + 3
1=1


t = 4
(4) - 2/(4) = 2/4 + 3
3.5 = 3.5